40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Nov 7 08:00:11 2016, Steve, Update, SUS, SRM -PRM sat. amp swap SRM.jpgSRM-UR_OK.png
    Reply  Mon Nov 7 16:05:55 2016, gautam, Update, SUS, PRM Sat. Box. Debugging D961289-B2.pdfPRMSatBoxtest.png
       Reply  Tue Nov 8 11:54:38 2016, gautam, Update, SUS, PRM Sat. Box. looks to be fixed 
Message ID: 12602     Entry time: Mon Nov 7 16:05:55 2016     In reply to: 12601     Reply to this: 12606
Author: gautam 
Type: Update 
Category: SUS 
Subject: PRM Sat. Box. Debugging 

Short summary of my Sat. Box. debugging activities over the last few days. Recall that the SRM Sat. Box has been plugged into the PRM suspension for a while now, while the SRM has just been hanging out with no electrical connections to its OSEMs.

As Steve mentioned, I had plugged in Ben's extremely useful tester box (I have added these to the 40m Electronics document sub-tree on the DCC) into the PRM Sat. Box and connected it to the CDS system over the weekend for observation. The problematic channel is LR.  Judging by Steve's 2 day summary plots, LR looks fine. There is some unexplained behavior in the UR channel - but this is different from the glitchy behaviour we have seen in the LR channel in the past. Moreover, subsequent debugging activities did not suggest anything obviously wrong with this channel. So no changes were made to UR. I then pulled out the PRM sat.box for further diagnostics, and also, for comparison, the SRM sat. box which has been hooked up to the PRM suspension as we know this has been working without any issues. 

Tracing out the voltages through the LED current driver circuit for the individual channels, and comparing the performance between PRM and SRM sat. boxes, I narrowed the problem down to a fault in either the LT1125CSW Quad Op-Amp IC or the LM6321M current driver IC in the LR channel. Specifically, I suspected the output of U3A (see Attachment #1) to be saturated, while all the other channels were fine. Looking at the spectrum at various points in the circuit with an SR785, I could not find significant difference between channels, or indeed, between the PRM/SRM boxes (up to 100kHz). So I decided to swap out both these ICs. Just replacing the OpAmp IC did not have any effect on the performance. But after swapping out the current buffer as well, the outputs of U3A and U11 matched those of the other channels. It is not clear to me what the mode of failure was, or if the problem is really fixed. I also checked to make sure that it was indeed the ICs that had failed, and not the various resistors/capacitors in the signal path. I have plugged in the PRM sat. box + tester box setup back into our CDS data acquisition for observation over a couple of days, but hopefully this does the job... I will update further details over the coming days.

I have restored control to PRM suspensions via the working SRM sat. box. The PRM Sat. Box and tester box are sitting near the BS/PRM chamber in the same configuration as Steve posted in his earlier elog for further diagnostics...


GV Edit 2230 hrs 7Nov2016: The signs from the last 6 hours has been good - see the attached minute trend plot. Usually, the glitches tend to show up in this sort of time frame. I am not quite ready to call the problem solved just yet, but I have restored the connections to the SRM suspension (the PRM and SRM Sat. Boxes are still switched). I've also briefly checked the SRM alignment, and am able to lock the DRMI, but the lock doesn't hold for more than a few seconds. I am leaving further investigations for tomorrow, let's see how the Sat. Box does overnight.

Attachment 1: D961289-B2.pdf  160 kB  | Hide | Hide all
D961289-B2.pdf
Attachment 2: PRMSatBoxtest.png  39 kB  Uploaded Mon Nov 7 22:24:37 2016  | Hide | Hide all
PRMSatBoxtest.png
ELOG V3.1.3-